OpenStudy (anonymous):

I am being asked to prove that the set { \[x \in \mathbb{R}: x^2 \le x\] } is equal to the interval [0,1]. Now I can see that any x on that interval this is true, but when they ask me to prove this, should I be using things like induction or contradiction? Or, is this provable just using the properties of an ordered field?

OpenStudy (anonymous):

\[x^2\leq x\iff x^2-x\leq 0\iff x(x-1)\leq 0\]

OpenStudy (anonymous):

i am not sure how fancy you need to be, but it is pretty obvious that if \(x<0\) we know \(x-1<0\) and so \(x(x-1)>0\)

OpenStudy (anonymous):

Ah right!

OpenStudy (anonymous):

And the same holds in the other direction for x > 1

OpenStudy (anonymous):

whereas if \(x>0, x-1<0\) i.e. if \(0<x<1\) we have \(x(x-1)<0\)

OpenStudy (anonymous):

yes it does

OpenStudy (anonymous):

Thanks a ton for this, for me it seems the hardest part of a proof is getting started. I appreciate the direction.

OpenStudy (anonymous):

yw